# OptiHive: Ensemble Selection for LLM-Based Optimization via Statistical Modeling

### v1

@@ Abstract
LLM-based solvers have emerged as a promising means of automating problem modeling and solving. However, they remain unreliable and often depend on iterative repair loops that result in significant latency. We introduce OptiHive, an LLM-based framework that produces high-quality solvers for optimization problems from natural-language descriptions without iterative self-correction. OptiHive uses a single batched LLM query to generate diverse components (solvers, problem instances, and validation tests) and filters out erroneous components to ensure fully interpretable outputs. Taking into account the imperfection of the generated components, we employ a statistical model to infer their true performance, enabling principled uncertainty quantification and solver selection. On tasks ranging from traditional optimization problems to challenging variants of the Multi-Depot Vehicle Routing Problem, OptiHive significantly outperforms baselines, increasing the optimality rate from 5% to 92% on the most complex problems.
@@ Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse domains. However, their application to complex optimization tasks remains hindered by unreliable code generation and self-evaluation. In practice, generated solvers often exhibit two distinct failure modes: hard computational errors (invalid syntax or runtime failure) that render code outright unusable, and soft quality deficiencies (incorrect algorithms, suboptimal solutions) that cannot be detected deterministically.
@@
Traditional LLM optimization pipelines either rely on iterative generate-evaluate-repair loops or fine-tuned models with integrated correction routines. These methods can address syntactic failures, but struggle to quantify the quality of solutions due to overconfidence in self-assessment. Further, they are subject to cyclical error patterns and suffer from high latency - that is, the time between receiving a problem description and returning a final solver. Test-based methods, which prompt the LLM to generate input-output pairs or simple verification functions using hard-coded values, can improve assessment in simple settings, but often fail to produce valid test cases for complex optimization problems, where ground-truth outputs cannot be obtained without solving the problem itself.
@@
In this work, we present OptiHive, a two-stage framework that decouples interpretability constraints from quality uncertainty, enabling both deterministic elimination of unusable components and statistical inference over the remaining ones. In the first stage, we simultaneously generate candidate solvers, problem instances, and validity tests. Interpretability, defined as the ability of a solver-instance-test triple to compile, execute, and return a syntactically valid result, is treated as a hard feasibility requirement. We formulate a mixed-integer linear program (MILP) to eliminate flawed components (solver, instances, and tests) and retain interpretable outputs only. The second stage then applies a latent-class model over the filtered elements, jointly estimating latent ground truth variables (instance feasibility and solution validity) as well as the error rates of solvers and tests.
@@
By quantifying the performance and trustworthiness of each solver, OptiHive provides principled uncertainty quantification and enables an informed solver selection. Unlike agent-based methods and multi-stage prompt pipelines that incrementally refine a single solution, our statistical ensemble method departs from the prevailing “generate-then-fix” paradigm: it simultaneously generates a set of candidate solvers, ranks them using their inferred true quality, and selects the most promising candidate. With minimal computational overhead from filtering, inference, and selection, OptiHive effectively serves two purposes: acting as a low-latency high performance LLM-based pipeline for solver generation, or as a wrapper around existing code synthesis frameworks to greatly improve performance through rigorous statistical inference instead of self-critique. The framework relies on the solver generation step to produce at least one correct solver. While deterministic sampling (e.g., temperature 0) may fail to yield any valid candidate, higher temperatures can introduce the diversity needed to uncover correct solvers. As a result, OptiHive can recover valid solutions even when deterministic generation fails.
@@
    Minimal and consistent latency through single batch LLM inference and parallelization. Our framework invokes the LLM exactly once to generate, in parallel, independent candidate solvers, problem instances, and validity tests, eliminating iterative self-correction loops. Combined with fully parallelizable cross-evaluation of solutions and tests, this design yields high-quality solvers with minimal latency.
@@
    Reliable solver selection via statistical inference. Previous work assumes that LLM self-evaluation provides trustworthy quality assessment, but recent research suggests that this assumption is fundamentally flawed due to LLMs’ poor self-critique abilities and systematic biases (Stechly, Valmeekam, and Kambhampati 2024; Chen et al. 2024a; Moon et al. 2025). Our framework addresses this by modeling all components (solvers, instances, and tests) as inherently noisy and employing rigorous statistical methods to estimate their true performance.
@@
Large language models (LLMs) have demonstrated remarkable abilities in reasoning, code generation, and problem solving across a wide range of domains (Brown et al. 2020; Achiam et al. 2023; Chen et al. 2021; Li et al. 2022), making them increasingly relevant tools for tackling complex computational tasks. Our work lies at the intersection of two streams of research within this field: LLMs for optimization problems and LLM-generated test functions.
@@
Chain-of-Thought (CoT) prompting (Wei et al. 2022) demonstrated that eliciting intermediate reasoning steps significantly improves performance on complex reasoning tasks in large-scale LLMs. Building on this, Composition of Experts (CoE) (Xiao et al. 2023) introduced an agent-based orchestrator that routes inputs to specialized LLM experts, improving scalability and customization in optimization pipelines. OptiMUS (AhmadiTeshnizi et al. 2024) and OptimAI (Thind et al. 2025) present end-to-end frameworks that automatically translate natural language problem descriptions into mathematical models solved by MILP solvers, validating the feasibility of large-scale optimization modeling. Both approaches incorporate self-correction loops, prompting the LLM to iteratively evaluate and revise its output until it passes a specified evaluation or reaches a maximum number of iterations. While this iterative refinement improves the correctness of generated models, it introduces significant latency and can be unreliable, particularly when the model repeatedly makes similar errors. Aligned with improvement loops, Optimization by PROmpting (OPRO)(Yang et al. 2023) proposes new candidates based on all previously generated solutions with their objective values, while Self-Guiding Exploration (SGE) (Iklassov et al. 2024) employs a thought-based framework that breaks each problem into simpler subtasks and then iteratively refines each “thought” through a fixed number of LLM calls. Drawing on evolutionary algorithms, Hercules (Wu et al. 2025) employs an outer loop that generates, evaluates, and selects heuristic candidates across successive generations. Parallel to these methods, LLMOPT (Jiang et al. 2024) and LLaMoCo (Ma et al. 2024) enhance the problem formulation and solver code generation via instruction-tuning. In contrast, our framework adopts a linear approach that eliminates self-correction loops by producing a set of independent candidate solvers, filtering out solvers returning non-interpretable results, and selecting the highest potential solver based on the parameters estimated by the latent-class model. This design achieves both low latency and high performance, overcoming a key bottleneck of existing LLM-driven optimization pipelines.
@@
Recent studies indicate that LLMs are poor and biased at self-critique (Huang et al. 2023; Stechly, Valmeekam, and Kambhampati 2024; Chen et al. 2024a; Moon et al. 2025). As a result, the automatic generation of external test functions, rather than relying on the model’s self-evaluation, has gained traction. Most works focus on unit test generation (Chen et al. 2022; Jain et al. 2024; Prasad et al. 2025; Chen et al. 2024b; Lin et al. 2025), where LLMs are prompted to write valid input-output pairs. These approaches significantly reduce manual testing effort and can achieve high code coverage, but are not amenable to complex optimization problems where the reasoning capabilities of the LLM are insufficient to produce valid input-output pairs. Only a handful of efforts have explored generating complete test functions (Yuan et al. 2023; Schäfer et al. 2023; Rao et al. 2023), but these remain restricted to simple assert-based checks over hard-coded inputs. In contrast, we generate reusable test functions that take input-output pairs and verify problem-specific invariants, such as constraint feasibility and objective value consistency. The decoupling of tests from specific inputs (i.e. problem instances) makes our framework thrifty, as each test can be reused to validate diverse solver-instance pairs.
@@
Figure 1 illustrates the two stages of OptiHive. In the first stage (generation of valid components), we efficiently produce executable components resulting in a set of interpretable outputs. Specifically, we invoke an LLM on a single batch inference to generate a pool of candidate solvers, problem instances, and validity tests. We compile and execute every solver-instance pair, and execute all tests on solver–instance pairs producing a solution. Then, we solve a MILP to remove a subset of solvers, instances, and tests, so that only interpretable outputs are retained. In the second stage (solver characterization and selection), we fit a latent-class model (Dawid and Skene 1979) on these interpretable outputs to jointly infer the feasibility of instances and each solver’s solution validity, and the type I/II error rates of both solvers and tests. Finally, we form a Pareto front of solvers based on their estimated error rates and expected objective values, and select a final solver either based on a priori methods or by minimizing a scalarized objective. The entire framework is given in Algorithm 1. We now describe each stage in detail, beginning with the generation of components.
@@
2: Perform single batch query to the LLM to obtain:
@@
8:  Execute solver $s$ on instance $i$ to obtain report $r_{s,i}$
@@
    Candidate Solvers $\bar{S}$. Each solver $s\in\bar{S}$ is function taking as input a problem instance $i$ and returning an optimization report with infeasible status if the instance admits no feasible solution, or the best solution found within the time limit at its corresponding objective value with either optimal or time_limit status.
@@
[…]
All prompts are provided in the Appendix. Importantly, every component in $\bar{S}$, $\bar{I}$, and $\bar{T}$ is independent of other components, so that all components can be generated simultaneously with a single batch of prompts. Note that no additional calls to the LLM are required during subsequent steps, which is key to the low latency of our framework.
@@
[…]
@@
[…]
@@
To retain a maximum number of components while filtering out all the non-interpretable triples, we formulate the following MILP:
@@
$S$$\triangleq\{s\in\bar{S}:x_{s}=1\}$
$I$$\triangleq\{i\in\bar{I}:y_{i}=1\}$
$T$$\triangleq\{t\in\bar{T}:z_{t}=1\}$
By construction, every triple in $S\times I\times T$ is interpretable. This MILP is tractable even with hundreds of solvers, instances, and tests, as the number of variables grows linearly with the number of components, and $\mathcal{U}$ exhibits a highly structured pattern since failure of a solver, instance, or test often induces multiple non-interpretable triples involving that component.
@@
$r_{s,i}$ =$=\left\{\begin{tabular}[]{cl}1&\text{if solver $s$ reports a solution}\\ 0&\text{otherwise.}\end{tabular}\right.$
@@
$r_{s,i,t}$ =$=\left\{\begin{tabular}[]{cl}1&\text{ if the solution of $(s,i)$ passes test $t$}\\ 0&\text{otherwise.}\end{tabular}\right.$
@@
$f_{i}$ =$=\left\{\begin{tabular}[]{cl}1&\text{if instance $i$ admits a feasible solution}\\ 0&\text{otherwise.}\end{tabular}\right.$
@@
$f_{s,i}$ =$=\left\{\begin{tabular}[]{cl}1&if \text{$r_{s,i}=1$ and the solution of $(s,i)$ is feasible}\\ 0&\text{otherwise.}\end{tabular}\right.$
@@
    $\mathbb{P}\left(r_{s,i}\;|\;f_{i}=0\right)=\alpha_{s}$
    $\mathbb{P}\left(r_{s,i}\;|\;f_{i}=1\right)=1-\beta_{s}$
    $\mathbb{P}\left(f_{s,i}\;|\;r_{s,i}=1,f_{i}=1\right)=\gamma_{s}$
    where $\alpha_{s}$, $\beta_{s}$, and $\gamma_{s}$ are false positive rate of solver, false negative rate, and rate of feasible reported solution, respectively, for solver $s$.
@@
    Tests. For each solver-instance pair with $r_{s,i}=1$, we aggregate the test outcomes as $C_{s,i}=\sum_{t\in T}r_{s,i,t}$. Test outcomes are not independent since some solutions are inherently harder to evaluate than others. However, conditioned on the latent variable $f_{s,i}$, the individual test results are assumed to be exchangeable. To capture this, we model $C_{s,i}$ Beta-Binomial distributions as:
    $C_{s,i}\;|\;f_{s,i}=0\sim\text{BetaBinomial}\left(T,a_{0},b_{0}\right)$
    $C_{s,i}\;|\;f_{s,i}=1\sim\text{BetaBinomial}\left(T,a_{1},b_{1}\right)$
To break the symmetry, we leverage the fact that tests rarely reject feasible solutions and use a strong prior $\text{Beta}(\bar{\alpha},\bar{\beta})$ with $\bar{\alpha}\gg\bar{\beta}$ on the tests true-positive rate $\frac{a_{1}}{a_{1}+b_{1}}$, which strongly favors values close to 1.
@@
$\theta_{k+1}=\operatorname*{arg\,max}_{\theta}\mathbb{E}_{\mathbf{F}\sim\mathbb{P}\left(\cdot\;|\;\mathbf{R},\theta_{k}\right)}\left[\ln\mathbb{P}\left(\mathbf{R},\mathbf{F}\;|\;\theta_{k}\right)\right]$ (2)
until convergence. The distribution of the latent variables $\{f_{i}\}_{i\in I}$, $\{f_{s,i}\}_{(s,i)\in S\times I}$ can then be estimated from $\theta^{\star}$. See Appendix B for details on update (2).
@@
$Z_{s}=\mathbb{E}_{i\sim\mathbb{P}\left(\cdot\;|\;r_{s,i}=1,f_{s,i}=1\right)}\left[z_{s,i}\right]$
@@
Specifically, $g(\theta^{\star},s)$ denotes the expected objective value of solver $s$ under penalties $P_{\text{miss}}$ when no solution is found for a feasible instance, and $P_{\text{fail}}$ when an infeasible solution is reported, respectively. By default, $P_{\text{miss}}$ and $P_{\text{fail}}$ can be set to $10\cdot Z_{\max}$, with $Z_{\text{max}}$ defined as the maximum absolute objective value reported across all solver-instance pairs. This ensures that both under-reporting and over-reporting solvers are severely penalized. This scalarized objective provides a basis for ranking solvers, from which the final solver is selected as $s^{\star}=\operatorname*{arg\,min}_{s\in S}g(\theta^{\star},s)$.
@@
Previous benchmark datasets for optimization from natural language, such as NLP4LP (AhmadiTeshnizi, Gao, and Udell 2023) and ComplexOR (Xiao et al. 2023), include problems that recent LLMs can now solve reliably, but where perfect scores remain difficult to achieve due to ambiguous or underspecified problem statements. For example, instances may describe integer-valued quantities while asking for LP formulations, leaving the model to infer the intended discretization. As a result, they do not allow for a reliable evaluation of solver performance, since failures may stem from prompt ambiguity rather than solver quality. To address this, we designed variants of the Multi-Depot Vehicle Routing Problem (MDVRP) and the Weighted Set Cover Problem (WSCP) that introduce increasing levels of complexity, enabling meaningful performance evaluation.
@@
We use OpenAI models (gpt-4.1-nano, gpt-4.1-mini, and o3), with the temperature fixed at the default value of 0.7. The specific model used is indicated for each problem. Runs are parallelized on an AMD EPYC 9734 processor. The number of expectation-maximization loops is limited to 100 iterations, and all individual runs complete in less than 1 second. This runtime is negligible compared to the time required for LLM inference or the overhead incurred by typical code synthesis pipelines.
Multi-Depots Vehicle Routing Problems
We study two variants of MDVRP (Toth and Vigo 2002):
@@
In the DCMDVRP, each vehicle’s total travel distance is constrained by a fixed upper bound. While this variant is less common than the standard MDVRP, it can typically be addressed by adding a single distance constraint per vehicle. Aside from rare formulation flaws, such as improperly allowing vehicle to reset their cumulated distance when visiting the depot of another vehicle, the formulation of this variant is conceptually very similar to MDVRP. In contrast, the MDVRP+OBS introduces significant geometric complexity: the 2D space contains line-segment obstacles that vehicles cannot cross, fundamentally altering the feasible routing space. Solvers that correctly address this problem typically follow a structured yet non-trivial pipeline: (i) augmenting the node set with obstacle endpoints, (ii) computing shortest paths via a visibility graph that excludes edges intersecting obstacles, (iii) solving a standard MDVRP using the resulting distance matrix, and (iv) mapping the MILP solution back to explicit routes. This process yields long, intricate code prone to oversimplification (e.g., by removing invalid edges instead of properly computing obstacle-aware shortest paths), which often leads to crashing, invalid, and suboptimal solutions.
@@
Figure 2: Feasibility and optimality rates on DCMDVRP and MDVRP+OBS across varying numbers of generated solvers, instances, and tests. The purple curve shows the optimality rate under perfect selection i.e. the probability that at least one of the generated solvers is optimal.
@@
Finally, the number of tests yields the smallest improvement in performance. Evaluating solution correctness is generally easier than generating a solution, and tests tend to be more reliable. As a result, the performance saturates quickly as more tests are added. In our experiments, we observe marginal improvement beyond five generated tests: once all failure modes are covered, additional tests contribute little to no new signal. While test diversity still matters to avoid blind spots and guard against rare flawed tests, we note that the cost–benefit curve for adding tests is substantially flatter than for increasing the number of solvers or instances in this experiment.
@@
The WSCP (Chvatal 1979; Caprara, Toth, and Fischetti 2000) can be illustrated through a practical scenario involving emitters and clients. Each emitter is characterized by a location, radius, and activation cost. An emitter is said to cover a client if the client lies within its coverage range. The objective is to select a minimum-cost subset of emitters so every client is covered by at least one active emitter. We consider three variants:
@@
The probabilistic variant introduces uncertain reliability of emitters. Here, each emitter $i$ fails independently with known probability $p_{i}$, and each client $j$ demands a minimum coverage probability $\pi_{j}$. Namely, if $x_{i}$ is a binary variables representing whether emitter $i$ is active, the solution must satisfy:

### v2

@@ Abstract
LLM-based solvers have emerged as a promising means of automating problem modeling and solving. However, they remain unreliable and often depend on iterative repair loops that result in significant latency. We introduce OptiHive, a framework that enhances any solver-generation pipeline to produce higher-quality solvers from natural-language descriptions of optimization problems. OptiHive uses a single batched generation to produce diverse components (solvers, problem instances, and validation tests) and filters out erroneous components to ensure fully interpretable outputs. Accounting for the imperfection of the generated components, we employ a statistical model to infer their true performance, enabling principled uncertainty quantification and solver selection. On tasks ranging from traditional optimization problems to challenging variants of the Multi-Depot Vehicle Routing Problem, OptiHive significantly outperforms baselines, increasing the optimality rate from 5% to 92% on the most complex problems.
Codehttps://github.com/mbscry/OptiHive
Extended versionhttps://arxiv.org/abs/2508.02503
@@ Introduction
Large language models (LLMs) have demonstrated remarkable abilities in reasoning, code generation, and problem solving across a wide range of domains (brown2020language; achiam2023gpt; chen2021evaluating; li2022competition), making them increasingly relevant tools for tackling complex computational tasks. However, their application to complex optimization tasks remains hindered by unreliable code generation and self-evaluation. In practice, generated solvers often exhibit two distinct failure modes: hard computational errors (invalid syntax or runtime failure) that render code outright unusable, and soft quality deficiencies (incorrect algorithms, suboptimal solutions) that cannot be detected deterministically.
@@
Traditional LLM optimization pipelines rely on iterative generate-evaluate-repair loops or fine-tuned models with integrated correction routines. These methods can address syntactic failures, but struggle to quantify the quality of solutions due to overconfidence in self-assessment. Moreover, the best-performing pipelines tend to involve long and often unpredictable latency, that is, the time between receiving a problem description and returning a final solver. Test-based methods, which prompt an LLM to generate input-output pairs or simple verification functions using hard-coded values, can improve assessment in simple settings, but often fail to produce valid test cases for complex optimization problems, where ground-truth outputs cannot be obtained without solving the problem itself.
@@
In this work, we present OptiHive, a two-stage framework that enhances any solver-generation pipeline by decoupling interpretability constraints from quality uncertainty, enabling both deterministic elimination of unusable components and statistical inference over the remaining ones. In the first stage, we simultaneously generate multiple candidate solvers, problem instances, and validity tests. Interpretability, defined as the ability of a solver-instance-test triple to compile, execute, and return a syntactically valid result, is treated as a hard feasibility requirement. We formulate an integer linear program (ILP) to eliminate flawed components (solvers, instances, and tests) and retain interpretable outputs only. The second stage then applies a latent-class model over the filtered elements, jointly estimating latent ground truth variables (instance feasibility and solution validity) as well as the error rates of solvers and tests. The framework is agnostic to the optimization class and can be applied to linear, convex, and nonconvex problems, including continuous and mixed-integer formulations.
@@
By quantifying the performance and trustworthiness of each solver, OptiHive provides principled uncertainty estimation and enables informed solver selection. Unlike agent-based methods and multi-stage prompt pipelines that incrementally refine a single solution, our statistical ensemble method departs from the prevailing “generate-then-fix” paradigm: it generates in parallel a set of candidate solvers from a given solver-generation pipeline, ranks the candidates by their inferred true quality, and selects the most promising one. OptiHive serves as a wrapper around existing solver-generation pipelines, substantially improving performance through rigorous statistical inference rather than self-critique. With minimal computational burden from filtering, inference, and selection, OptiHive adds negligible latency to the pipeline it wraps, and enables high performance with minimal overhead when paired with lightweight generation procedures such as single-shot LLM inference.
The framework relies on the solver generation step to produce at least one correct solver. While deterministic sampling (e.g., temperature 0) may fail to yield any valid candidate, higher temperatures can introduce the diversity needed to uncover correct solvers. As a result, OptiHive can recover valid solutions even when deterministic generation fails.
@@
    Minimal and consistent latency through single batched inference and parallelization. Our framework generates, in a single parallel step, independent candidate solvers, problem instances, and validity tests, eliminating iterative self-correction loops. Combined with fully parallelizable cross-evaluation of solutions and tests, this design yields high-quality solvers with minimal latency.
@@
    Reliable solver selection via statistical inference. Previous work assumes that LLM self-evaluation provides trustworthy quality assessment, but recent research suggests that this assumption is fundamentally flawed due to LLMs’ poor self-critique abilities and systematic biases (stechly2024self; chen2024magicore; moon2025don). Our framework addresses this by modeling all components (solvers, instances, and tests) as inherently noisy and employing rigorous statistical methods to estimate their true performance.
@@
Our work lies at the intersection of two streams of research within this field: LLMs for optimization problems and LLM-generated test functions.
@@
Chain-of-Thought (CoT) prompting (wei2022chain) demonstrated that eliciting intermediate reasoning steps significantly improves performance on complex reasoning tasks in large-scale LLMs. Building on this, Composition of Experts (CoE) (CoE) introduced an agent-based orchestrator that routes inputs to specialized LLM experts, improving scalability and customization in optimization pipelines. OptiMUS (OptiMUS) and OptimAI (OptimAI) present end-to-end frameworks that automatically translate natural language problem descriptions into mathematical models, validating the feasibility of large-scale optimization modeling. Both approaches incorporate self-correction loops, prompting the LLM to iteratively evaluate and revise its output until it passes a specified evaluation or reaches a maximum number of iterations. While this iterative refinement improves the correctness of generated models, it introduces significant latency and can be unreliable, particularly when the model repeatedly makes similar errors. Aligned with improvement loops, Optimization by PROmpting (OPRO) (yang2023large) proposes new candidates based on all previously generated solutions with their objective values, while Self-Guiding Exploration (SGE) (iklassov2024self) employs a thought-based framework that breaks each problem into simpler subtasks and then iteratively refines each “thought” through a fixed number of LLM calls. Drawing on evolutionary algorithms, Hercules (wu2025efficient) employs an outer loop that generates, evaluates, and selects heuristic candidates across successive generations. Parallel to these methods, LLMOPT (LLMOPT) and LLaMoCo (Llamoco) enhance the problem formulation and solver code generation via instruction-tuning. In contrast, our framework adopts a linear approach that eliminates self-correction loops by producing a set of independent candidate solvers, filtering out solvers returning non-interpretable results, and selecting the highest potential solver based on the parameters estimated by the latent-class model. This design achieves both low latency and high performance, overcoming a key bottleneck of existing LLM-driven optimization pipelines.
@@
Recent studies indicate that LLMs are poor and biased in self-critique (huang2023large; stechly2024self; chen2024magicore; moon2025don). As a result, the automatic generation of external test functions, rather than relying on the model’s self-evaluation, has gained traction. Most works focus on unit test generation (chen2022codet; jain2024livecodebench; prasad2025learning; chen2024chatunitest; lin2025learning), where LLMs are prompted to write valid input-output pairs. These approaches significantly reduce manual testing effort and can achieve high code coverage, but are not amenable to complex optimization problems where the reasoning capabilities of the LLM are insufficient to produce valid input-output pairs. Only a handful of efforts have explored generating complete test functions (yuan2023no; schafer2023empirical; rao2023cat), but these remain restricted to simple assert-based checks over hard-coded inputs. In contrast, we generate reusable test functions that take input-output pairs and verify problem-specific invariants, such as constraint feasibility and objective value consistency. The decoupling of tests from specific inputs (i.e. problem instances) makes our framework thrifty, as each test can be reused to validate diverse solver-instance pairs.
@@
Figure 1 illustrates the two stages of OptiHive. In the first stage (generation of valid components), we efficiently produce executable components, resulting in a set of interpretable outputs. Specifically, we perform a single batch generation to obtain a pool of candidate solvers, problem instances, and validity tests. We compile and execute every solver-instance pair, and execute all tests on solver–instance pairs producing a solution. Then, we solve an ILP to remove a subset of solvers, instances, and tests, so that only interpretable outputs are retained. In the second stage (solver characterization and selection), we fit a latent-class model (dawid1979maximum) on these interpretable outputs to jointly infer the feasibility of instances and each solver’s solution validity, and the type I/II error rates of both solvers and tests. Finally, we rank solvers with a scalarized objective that integrates their estimated error rates and expected objective values, and select the top-ranked candidate. The entire framework is given in Algorithm 1. We now describe each stage in detail, beginning with the generation of components.
@@
2: Perform single batch generation to obtain:
@@
8:  Execute solver $s$ on instance $i$ to obtain $r_{s,i}$
@@
    Candidate Solvers $\bar{S}$. Each solver $s\in\bar{S}$ is a function taking as input a problem instance $i$ and returning an optimization report with infeasible status if the instance admits no feasible solution, or the best solution found within the time limit at its corresponding objective value with either optimal or time_limit status.
@@
    Validity Tests $\bar{T}$. Each test $t\in\bar{T}$ is a function taking as input an instance-solution pair $(i,x)$ and returning True if and only if the solution $x$ is feasible for $i$ and its true objective value matches the reported objective value.
All prompts are provided in Appendix A of the extended version of this paper. Importantly, every component in $\bar{S}$, $\bar{I}$, and $\bar{T}$ is independent of other components, so that all components can be generated simultaneously. The set $\bar{S}$ can be obtained using a more refined generation pipeline. In any case, no additional calls to the LLM are required during subsequent steps, which is key to the low latency of our framework.
@@
Note that all LLM-generated components are subject to being syntactically or semantically invalid. For each tuple $(s,i)\in\bar{S}\times\bar{I}$, we attempt to obtain an optimization report by running $s$ on $i$. We say that the tuple $(s,i)$ is interpretable if $s$ and $i$ compile, running $s$ on $i$ does not raise an error during execution, and the report of $(s,i)$ contains a status field with an interpretable value. For every test $t$, we say that the triple $(s,i,t)$ is interpretable when test $t$ compiles and either:
@@
    $(s,i)$ is interpretable with a report containing a solution, running test $t$ on the solution of $(s,i)$ does not raise error during execution, and returns a boolean.
@@
To retain a maximum number of components while filtering out all the non-interpretable triples, we formulate the following ILP:
@@
$S$ $\triangleq\{s\in\bar{S}:x^{\star}_{s}=1\}$
$I$ $\triangleq\{i\in\bar{I}:y^{\star}_{i}=1\}$
$T$ $\triangleq\{t\in\bar{T}:z^{\star}_{t}=1\}$
By construction, every triple in $S\times I\times T$ is interpretable. This ILP is tractable even with hundreds of solvers, instances, and tests, as the number of variables grows linearly with the number of components, and $\mathcal{U}$ exhibits a highly structured pattern since failure of a solver, instance, or test often induces multiple non-interpretable triples involving that component.
@@
$r_{s,i}$ $=\left\{\begin{tabular}[]{cl}1&\text{if solver $s$ reports a solution}\\ 0&\text{otherwise.}\end{tabular}\right$.
@@
$r_{s,i,t}$ $=\left\{\begin{tabular}[]{cl}1&\text{ if the solution of $(s,i)$ passes test $t$}\\ 0&\text{otherwise.}\end{tabular}\right$.
@@
$f_{i}$ $=\left\{\begin{tabular}[]{cl}1&\text{if instance $i$ admits a feasible solution}\\ 0&\text{otherwise.}\end{tabular}\right$.
@@
[…]
@@
    $\mathbb{P}\left(r_{s,i}=1\mid f_{i}=0\right)=\alpha_{s}$
    $\mathbb{P}\left(r_{s,i}=1\mid f_{i}=1\right)=1-\beta_{s}$
    $\mathbb{P}\left(f_{s,i}=1\mid r_{s,i}=1,f_{i}=1\right)=\gamma_{s}$
    where $\alpha_{s}$, $\beta_{s}$, and $\gamma_{s}$ are false positive rate, false negative rate, and rate of feasible solutions among reports on positive instances, respectively, for solver $s$.
@@
    Tests. For each solver-instance pair with $r_{s,i}=1$, we aggregate the test outcomes as $C_{s,i}=\sum_{t\in T}r_{s,i,t}$. Test outcomes are not independent since some solutions are inherently harder to evaluate than others. However, conditioned on the latent variable $f_{s,i}$, the individual test results are assumed to be exchangeable. To capture this, we model $C_{s,i}$ with Beta-Binomial distributions as:
    $C_{s,i}\mid f_{s,i}=0\sim\text{BetaBinomial}\left(|T|,a_{0},b_{0}\right)$
    $C_{s,i}\mid f_{s,i}=1\sim\text{BetaBinomial}\left(|T|,a_{1},b_{1}\right)$
To break the symmetry, we leverage the fact that tests rarely reject feasible solutions and use a strong prior $\text{Beta}(\bar{\alpha},\bar{\beta})$ with $\bar{\alpha}\gg\bar{\beta}$ on the tests’ true-positive rate $\frac{a_{1}}{a_{1}+b_{1}}$, which strongly favors values close to 1.
@@
$\theta_{k+1}=\operatorname*{arg\,max}_{\theta}\mathbb{E}_{\mathbf{F}\sim\mathbb{P}\left(\cdot\mid\mathbf{R},\theta_{k}\right)}\left[\ln\mathbb{P}\left(\mathbf{R},\mathbf{F}\mid\theta\right)\right]$ (2)
until convergence. The distribution of the latent variables $\{f_{i}\}_{i\in I}$, $\{f_{s,i}\}_{(s,i)\in S\times I}$ can then be estimated from $\theta^{\star}$. See Appendix B in the extended version for details on update (2).
@@
$Z_{s}=\mathbb{E}_{i\sim\mathbb{P}\left(\cdot\mid r_{s,i}=1,f_{s,i}=1\right)}\left[z_{s,i}\right]$
@@
Specifically, $g(\theta^{\star},s)$ denotes the expected objective value of solver $s$ under penalties $P_{\text{miss}}$ when no solution is found for a feasible instance, and $P_{\text{fail}}$ when an infeasible solution is reported, respectively. By default, $P_{\text{miss}}$ and $P_{\text{fail}}$ can be set to $10\cdot Z_{\text{max}}$, with $Z_{\text{max}}$ defined as the maximum absolute objective value reported across all solver-instance pairs. This ensures that both under-reporting and over-reporting solvers are severely penalized. This scalarized objective provides a basis for ranking solvers, from which the final solver is selected as $s^{\star}=\operatorname*{arg\,min}_{s\in S}g(\theta^{\star},s)$.
@@
Previous benchmark datasets for optimization from natural language, such as NLP4LP (ahmaditeshnizi2023optimus) and ComplexOR (CoE), include problems that recent LLMs can now solve reliably, but where perfect scores remain difficult to achieve due to ambiguous or underspecified problem statements. For example, instances may describe integer-valued quantities while asking for LP formulations, leaving the model to infer the intended discretization. As a result, they do not allow for a reliable evaluation of solver performance, since failures may stem from prompt ambiguity rather than solver quality. To address this, we designed variants of the Multi-Depot Vehicle Routing Problem (MDVRP) and the Weighted Set Cover Problem (WSCP) that introduce increasing levels of complexity, enabling meaningful performance evaluation.
@@
We use OpenAI models (gpt-4.1-nano, gpt-4.1-mini, and o3), with the temperature fixed at the default value of 0.7. The specific model used is indicated for each problem. Runs are parallelized on an AMD EPYC 9734 processor. The number of expectation-maximization loops is limited to 100 iterations, and all individual runs complete in less than 1 second. This runtime is negligible compared to the time required for LLM inference or the overhead incurred by typical code generation pipelines.
Multi-Depot Vehicle Routing Problems
We study two variants of MDVRP (toth2002vehicle):
@@
In the DCMDVRP, each vehicle’s total travel distance is constrained by a fixed upper bound. While this variant is less common than the standard MDVRP, it can typically be addressed by adding a single distance constraint per vehicle. Aside from rare formulation flaws, such as improperly allowing vehicles to reset their cumulated distance when visiting the depot of another vehicle, the formulation of this variant is conceptually very similar to MDVRP. In contrast, the MDVRP+OBS introduces significant geometric complexity: the 2D space contains line-segment obstacles that vehicles cannot cross, fundamentally altering the feasible routing space. Solvers that correctly address this problem typically follow a structured yet non-trivial pipeline: (i) augmenting the node set with obstacle endpoints, (ii) computing shortest paths via a visibility graph that excludes edges intersecting obstacles, (iii) solving a standard MDVRP using the resulting distance matrix, and (iv) mapping the ILP solution back to explicit routes. This process yields long, intricate code prone to oversimplification (e.g., by removing invalid edges instead of properly computing obstacle-aware shortest paths), which often leads to crashing, invalid, and suboptimal solutions.
@@
Figure 2: Feasibility and optimality rates on DCMDVRP and MDVRP+OBS across varying numbers of generated solvers, instances, and tests. The purple curve shows the rate under perfect selection i.e. the probability that at least one of the generated solvers is optimal.
@@
Finally, the number of tests yields the smallest improvement in performance. Evaluating solution correctness is generally easier than generating a solution, and tests tend to be more reliable. As a result, the performance saturates quickly as more tests are added. In our experiments, we observe marginal improvement beyond five generated tests: once all failure modes are covered, additional tests add little value. While test diversity still matters to avoid blind spots and guard against rare flawed tests, we note that the cost–benefit curve for adding tests is substantially flatter than for increasing the number of solvers or instances in this experiment.
@@
The WSCP (chvatal1979greedy; caprara2000algorithms) can be illustrated through a practical scenario involving emitters and clients. Each emitter is characterized by a location, radius, and activation cost. An emitter is said to cover a client if the client lies within its coverage range. The objective is to select a minimum-cost subset of emitters so every client is covered by at least one active emitter. We consider three variants:
@@
The probabilistic variant introduces uncertain reliability of emitters. Here, each emitter $i$ fails independently with known probability $p_{i}$, and each client $j$ demands a minimum coverage probability $\pi_{j}$. Namely, if $x_{i}$ is a binary variable representing whether emitter $i$ is active, the solution must satisfy: